OpenStudy (anonymous):

The sum of digits of a three-digit number is 16. If the hundreds' digit is half the units' digit and equal to the tens' digit, what is the number?

OpenStudy (anonymous):

So you have a number of the form xyz, and you're told that \(x+y+z=16\). What else are you told about these digits?

OpenStudy (anonymous):

Those are the only information given. I guess then that x= 1/2z and y=x.

OpenStudy (anonymous):

Exactly! So in the equation \(x+y+z=16\), you can replace the y, and the z with terms involving x. What does that look like?

OpenStudy (anonymous):

Ooooh I get it. Umm, x+x+2x=16 4x=16 x=4 :)

OpenStudy (anonymous):

congrats! Now you have to figure out what y, and z are.

OpenStudy (anonymous):

Lol. x=4 then because y=x, y=4. then x=1/2z or, 2x=z. So, 2(4)=z which is 8. The number's 484. Thanks :D

OpenStudy (anonymous):

Close! It's 448, remember that it's xyz, not xzy :)

OpenStudy (anonymous):

Lol. My bad. :)